OpenStudy (anonymous):

3/x+1 + 4x/x-2 Explain please?

OpenStudy (anonymous):

can you rtype this? it doesn't make since

OpenStudy (anonymous):

explain what? it's an expression involving a fractions of polynomials?

OpenStudy (anonymous):

\[\frac{ 4x }{ x-2 }+\frac{ 3 }{ x+1 }\] How am I supposed to do it??

OpenStudy (anonymous):

well to add fractions of polynomials they must have the same divisor. So if you have: a/b+c/d you do ad/bd+bc/bd=(ad+bc)/bd, that is assuming the polynomials b and have no common factors. so in your case you have (4x(x+1)+3(x-2))/((x-2)(x+1)